# Decision Theoretic Foundations for Experiments Evaluating Human Decisions

### v3

@@ Abstract.
How well people use information displays to make decisions is of primary interest in human-centered AI, model explainability, data visualization, and related areas. However, what constitutes a decision problem, and what is required for a study to establish that human decisions could be improved remain open to speculation. We propose a widely applicable definition of a decision problem synthesized from statistical decision theory and information economics as a standard for establishing when human decisions can be improved in HCI. We argue that to attribute loss in human performance to forms of bias, an experiment must provide participants with the information that a rational agent would need to identify the utility-maximizing decision. As a demonstration, we evaluate the extent to which recent evaluations of decision-making from the literature on AI-assisted decisions achieve this criteria. We find that only 10 (26%) of 39 studies that claim to identify biased behavior present participants with sufficient information to characterize their behavior as deviating from good decision-making in at least one treatment condition. We motivate the value of studying well-defined decision problems by describing a characterization of performance losses they allow us to conceive. In contrast, the ambiguities of a poorly communicated decision problem preclude normative interpretation. We conclude with recommendations for practice.
@@ 1. Introduction
Providing informative interfaces to aid decision-making is a common goal in research in human-centered artificial intelligence (HCAI), visualization, and related areas. For example, visualization researchers emphasize assisting decision-making as a vital goal of data visualization (Dimara and Stasko, 2021). Similarly, in human-centered AI, a proliferation of empirical studies on human decision behavior have been deemed “necessary to evaluate the effectiveness of AI technologies in assisting decision making, but also to form a foundational understanding of how people interact with AI to make decisions” (Lai et al., 2023).
@@
Research areas in HCI and related fields that rely heavily on decision-making lack a consensus on what a decision is, and how decisions are a distinct form of human behavior compared to tasks that are not considered decisions. Human decisions from prediction and other information displays are often evaluated in studies with the goal of identifying forms of bias: ways in which study participants’ behavior deviates from expectations of good decision-making. However, reviewing the range of tasks that get filed under decision-making, it is difficult to ascertain what common task structure makes these studies about decision-making, much less what properties a study needs to have in order for it to be capable of distinguishing human decisions as flawed.
@@
As examples of the inconsistency and underdefinition of good decision-making in the HCI literature, colloquial definitions describe a decision task as a choice between alternatives, or a choice where the stakes are high (such as deciding which of multiple hurricane forecasts is more accurate (Padilla et al., 2015), or whether to screen child welfare cases (Zytek et al., 2021)), neither of which isolates a clear set of assumptions that must hold for a judgment to qualify as a decision. Other accounts cast decision-making as a higher-level task akin to sensemaking (Dimara and Stasko, 2021). For some, the point of studying human decisions is to compare them to a definition of idealized behavior. However, sometimes researchers define this standard in a way that is common across studies (e.g., a better forecast comes closer to the true probability of the event (Green and Chen, 2019; Fogliato et al., 2021)) while other times different definitions of the same concept are used (e.g., appropriate reliance as agreeing with an AI when it is correct and disagreeing when it’s not (Bansal et al., 2021; Buçinca et al., 2021) versus providing a higher subjective rating of trust when the model is correct (Bussone et al., 2015) versus choosing the better of the human versus the AI prediction to bet on for a future task (Dietvorst et al., 2015)). Sometimes the point is to study seemingly “subjective decisions,” like how people react to music recommendations (Kulesza et al., 2012), or decide whether a house seems to match its valuation (Prabhudesai et al., 2023), where it can seem unclear whether normative interpretations of task performance can be applied. Not only is a more coherent definition of decision needed, as recent survey papers on decision making in visualization and human-centered AI call for (Dimara and Stasko, 2021; Lai et al., 2021), but such a definition must clearly define requirements for identifying forms of bias or performance loss in human decision-making.
We argue for the identification of a minimum set of theoretical commitments for the task induced by a study of human decision-making from information displays to be well-defined. By well-defined we mean that it is possible to establish normative behavior, and consequently to conclude that observed human decisions are flawed in some way. We draw on statistical decision theory and expected utility theory (Savage, 1972; Steele and Stefánsson, 2015; Von Neumann and Morgenstern, 2007), which provide a rigorously derived and widely applicable framework for studying decision-making. Information economics formalizes information structures for decision problems(Bergemann and Morris, 2019), which may by induced by design choices about how to visualize or explain model predictions.
We contribute a widely applicable definition of a well-defined decision problem, and an argument for the value of this approach to HCAI and related research concerning decision-making from data-driven interfaces. Our first contribution is to establish and motivate a minimum set of components that must be defined for a decision problem to allow identification of an optimal decision (or set of decisions) so as to identify bias in human decisions. Using the concept of a rational Bayesian agent, we show how performance loss can only be conceived of when a study participant could in theory identify how to optimize their behavior (i.e., identify the normative decision) from the information they are provided. We find that while 39 of a sample of 46 existing studies draw conclusions about shortcomings in human decisions with prediction displays, only for 10 (26%) of these is the study task well-defined in at least one treatment condition, in the sense that participants are given sufficient information to identify, at least in principle, the best decision. We use examples to illustrate the epistemological risks to such conclusions and provide recommendations for experimenters to improve the interpretability of their study results.
@@ 2. Defining a Decision Problem
Our definition of a decision problem and corresponding standard for optimal behavior is intended for controlled evaluation of human behavior, a.k.a. normative decision research. Behavioral data (which may be produced by humans or by simulation) is collected under controlled conditions, with the goal of learning about behavior induced by the provision of information. Such studies are frequently used to describe the quality of human performance in some situation (e.g., how well people make decisions from displays in strategic settings (Zhang et al., 2023)), to rank different assistive elements by human performance (e.g, different visualizations (Correll et al., 2018) or AI explanation strategies (Liu et al., 2021)), or to test a hypothesis about how humans make decisions or what will help them do better (e.g., cognitive forcing functions will improve AI-assisted decisions (Buçinca et al., 2021)). Critically, such evaluative studies require that ground truth can be identified for any states that the study participant is queried about.
@@ 2.1. Decision Problem
The data-generating model is a distribution over values of the uncertain state and signals that provide information to the agent about the state. The data-generating model associates the state with the information that the agent is presented with (the signal) in order to inform their decision. Since the signal is correlated with the state, if the participant understands it in the context of the decision problem, they can improve their performance compared to a case where they do not have access to the signal.
When we present a person (hereafter agent) with a decision problem, we give them a choice of response (action) from some set of possible responses (action space) . In decision tasks corresponding to prediction tasks, the action space is a probabilistic belief over the state space, i.e., . The agent faced with a decision task must consider the probability of different states induced by the signal, and choose the best action under the scoring rule, a function that assigns a score representing the quality of the action for the realized state: .
We refer to the process that produces signals that inform on the uncertain state as a signaling policy. The relevance of the signal to the decision task is captured by the correlation between the signal and the uncertain state. More formally, a signaling policy induces an information structure that is given by a joint distribution over signals and states . This joint distribution assigns to each realization a probability, denoted .
@@
Under this definition, the formalization of the information structure enables a definition of normative behavior. Summarizing information about , (as well as ) provides a basis for the experimenter to establish expectations under optimal use of the information.
@@ 2.2. Normative Behavior
2.2.1. Rational Belief Formation
To define the normative action for a decision problem, we must first characterize how we would ideally expect the agent to form the posterior beliefs about the state given the signal . A theory of optimal belief construction starts by assuming that the ideal agent perfectly perceives . Once perceived, how does inform the agent’s beliefs about ? Generally, we assume that the agent has some baseline set of beliefs about that they could use to choose an action in the absence of . Encountering leads the agent to Bayesian update their beliefs.
where is a normalizing factor. Note that the definition in Equation 3 implies that to calculate the agent knows :
Given these posterior beliefs, we use Equation 2 to identify the action that a fully rational agent would choose in order to maximize their expected utility under .
@@ 2.3. Example
To make these concepts concrete, imagine a study of AI-advised decisions in which people make recidivism decisions upon seeing features of a defendant. The state is whether the person will commit another crime, i.e., recidivate, within some duration after being released, which can take the value 0 for no or 1 for yes (). The action space presented to the agent is whether or not to release the defendant ().
The data-generating model is the joint distribution over the state (recidivate or not) and signals comprised of defendant feature values and AI prediction information. Assume that agents are presented with a signal in the form of a set of feature values describing a defendant (age, prior arrests, prior felonies, etc.) from which they must decide on an action. Imagine that the signal also includes the prediction from a model that agents are told was trained on a historical dataset of labeled examples, along with the probability that the model’s prediction is correct. Agents can optionally use this information to inform their decision. Here, the joint distribution assigns a probability, for each possible value of ({0, 1}), to each possible signal (a combination of feature values and model prediction with associated probability of correctness) used as a trial in the experiment.
If we did not show the agent any feature values at all (nor model predictions), we would expect them to make their choice based on their a priori sense of the probability that a randomly drawn defendant would recidivate. The prior distribution gives the probability that a defendant recidivates as defined by the data-generating model before viewing the signal given on an particular trial.
Experiment designs can vary in the amount of information they provide participants about . What is critical is that a participant has some opportunity to obtain information about . For example, imagine a version of the recidivism study in which the agent is given only the feature values and model prediction, with no information to link these to the state (i.e., no probability that the model prediction is correct, and no feedback on the realized state after their decision). Consequently, they have no opportunity to learn about , which is necessary to establish a normative decision.
2.4. Additional Remarks on Scoring Rules
Returning to the recidivism example, assume the researcher uses a binary scoring rule to map the ground truth realization of the state ( for not recidivate, recidivate) and the participant’s selected action from the action space (), say, releasing the defendant (), to an outcome. For example, the agent might be scored by the following rule, which imposes the greatest penalty on false positives (i.e., acting as if the defendant will recidivate when they will not):
@@ 2.4. Additional Remarks on Scoring Rules
Whether the researcher uses the same scoring rule for both incentives and evaluation versus two different rules has important implications for interpreting the results. For the recidivism example, if the researchers analyzed the collected responses using the above rule (i.e., reporting performance in payoff space) in addition to communicating it to participants for incentivization, then the utility associated with performance is defined by a single scoring rule, and the interpretation of a “good” decision is consistent across incentivization and evaluation, which is clearly beneficial for interpreting the results.
One way around this problem when using two different rules is to make sure that the rule used for incentivization has certain properties. Specifically, among scoring rules for beliefs, we can distinguish those that are proper scoring rules: scoring rules for which the optimal action is to predict the true distribution, i.e., . When a proper scoring rule is used, the researcher can assume that an agent who understands the rule properly will report their true beliefs, as it is not possible to score higher by reporting some alternative beliefs. These beliefs can then be plugged into a different scoring rule without the risk of misinterpretation.
Applying the construction in Equation 6 to a binary decision gives a threshold rule where for each state, all reported probabilities below a threshold receive one score and those above receive another score. Since the payoff depends only weakly on the belief, such scoring rules may be less sensitive to the agents’ understanding of the signal. For this reason, including a decision problem that directly also elicits beliefs using a proper-scoring rule is a good idea whenever eliciting responses using an action space that is not over beliefs, as it enables plugging the elicited beliefs into other scoring rules.111See, e.g., Wu et al. (2023)’s analysis of (Kale et al., 2020).
The choice of possible scoring rules depends on the nature of the action space. In general, a coarser action space (i.e., with fewer choices available to participants) will produce less informative responses for the purpose of resolving observed performance relative to a normative standard. If, for example, the researcher designed the recidivism decision study to also elicit the agent’s prediction of the probability of recidivism, each trial would provide more information, and they could employ a proper scoring rule such as squared loss (i.e., Brier or quadratic rule) or log score to score the accuracy of the probabilistic beliefs. Different proper scoring rules imply different expectations about the set of possible decisions (Levinstein, 2017).
2.5. Extensions
Learning and Feedback Given a definition of normative behavior dictated by Bayesian decision theory, we would expect the decision-maker to update their beliefs given any information that informs on the joint distribution between signals and states. One source of such information is showing the agent the realization of the state after they make their choice. Imagine that participants in the recidivism decision study see feedback (i.e., are told the realized state, whether or not the defendant recidivated) after each decision they make. Assuming the signal is not revealing of the posterior probability of the state (recidivate or not), we would naturally expect the participants to learn about from this feedback as they complete trials, affecting their posterior beliefs (Equations 3 and 4).
Decision-Making Under Risk and Certainty The phrase “decision-making under uncertainty” is defined in the economics literature as the case where the decision-maker is not directly given the probability distribution over the state but instead must infer it. Alternatively, decision-making under risk describes the situation when this distribution is assumed to be known to the decision-maker (Steele and Stefánsson, 2015). Under such an assumption we simply replace in Equation 2 with the known distribution.
We can also easily account for so-called decision-making under certainty, where the state is deterministic222For any uncertain decision, it is possible to devise an analogous set of judgments where states are deterministic and enumerate the entire state space, an approach which has been used to isolate the effect of uncertainty from the complexity of reasoning about multiple contingencies (Martínez-Marquina et al., 2019). Under certainty, we no longer need to compute expectation to calculate the optimal decision:
3. Defining Bias in Responses to a Decision Problem
The primary motivations for using the above framework are epistemological in nature: they concern what knowledge can be gained from the results of a study. Our primary claim in this work is that to interpret the results of an experiment on human decision-making as evidence that human decisions are flawed in some way, the experiment must provide the participant with enough information to in principle identify the normative decision that will be used to judge their behavior. In other words, does the experiment give the participant enough information to align their understanding of the decision problem with its normative interpretation?
When an experiment does not provide sufficient information for the participant to solve the task in principle, there is the potential for multiplicity concerning the data-generating model. Such multiplicity has implications for interpreting the study results. As an example, consider the common approach to defining appropriate reliance on AI as agreeing with the AI when it is correct and disagreeing when it is not (Bansal et al., 2021; Gonzalez et al., 2020; Wang and Yin, 2021; Yang et al., 2020). Imagine a user receives information about the model’s overall performance unconditional on the predicted class. Consider two data-generating models that are consistent with the provided information: one where errors are always false positives, one where errors are always false negatives. Imagine a user who can only correct false positives. If we observe the user disagreeing with the model prediction some observed proportion of the time, and getting some of those instances wrong we would say they are overrelying under the first data-generating model. If they disagree with some model predictions under the second data-generating model, we would say they are underrelying. However, both are compatible with the information they were given.
3.1. Scope
It is worth making several remarks on the scope of our argument. It is important to recognize that even if the researchers do not think about the elements of their experiment in terms of the components identified above (i.e., statespace, data-generating model, signaling policy, action space, scoring rule), the design of their study still induces a decision problem whenever two simple conditions hold (Wu et al., 2023). First, there must be some uncertainty about the state (i.e., there is more than one data scenario used to generate stimuli)333A single trial study where all participants view the same signal would not qualify., and second, the researchers believe it is possible to provide a scoring rule that is consistent across participants (eliminating, for example, studies of emotional responses).
It should also be noted that our argument is that the interpretation of results is restricted when a well-defined decision problem is not conveyed to participants. Researchers can certainly conduct and report on studies that do not conform to this approach, but they cannot conclude that people made bad decisions if they were not given sufficient information to know what a good decision is. For example, imagine a study comparing the decision performance of users with and without the use of an AI, where no prior is given to participants, nor a scoring rule. Imagine that the performance measured in terms of accuracy is higher access to an AI. Given an appropriately designed randomized experiment the researcher could conclude that there is a causal effect of the AI on decisions. However, they could not claim that people are making worse or biased decisions without the AI, because they would be scoring their decisions using information they never provided to participants. When we say that one condition did better, we should acknowledge that we did not give them a chance to fully understand and optimize their behavior for the problem that we analyzed.
Finally, a common misinterpretation of the use of a rational agent model is that we are implying that human behavior is expected to be perfectly rational, or approximately rational. This is not the case. While a rational agent approach formally justifies what it means to best use information to inform the selection of an action using statistical decision theory, participants in our experiment may use different strategies. The point is that without a well-defined normative response dictating what it means to best respond to the information we have provided, we are hard-pressed to evaluate participant behavior. Even if we do not expect this procedure to characterize what human agents do, it provides a basis for interpreting our results via deviations from a well-understood standard.
We we describe below, when the conditions above are met, we can characterize forms of bias as loss in decision performance according to four possible sources.

### v5

@@ Abstract.
Decision-making with information displays is a key focus of research in areas like human-AI collaboration and data visualization. However, what constitutes a decision problem, and what is required for an experiment to conclude that decisions are flawed, remain open to speculation. We present a widely applicable definition of a decision problem synthesized from statistical decision theory and information economics. We claim that to attribute loss in human performance to bias, an experiment must provide the information that a rational agent would need to identify the normative decision. We evaluate whether recent empirical research on AI-assisted decisions achieves this. We find that only 10 (26%) of 39 studies that claim to identify biased behavior present participants with sufficient information to make this claim in at least one treatment condition. We motivate the value of studying well-defined decision problems by describing a characterization of performance losses they allow us to conceive.
@@ 1. Introduction
Identifying effective, informative interfaces to aid decision-making is a common goal in research in human-centered artificial intelligence (HCAI), visualization, and related areas. For example, visualization researchers emphasize assisting decision-making as a vital goal of data visualization (Dimara and Stasko, 2021). Similarly, in human-centered AI, a proliferation of empirical studies on human decision behavior have been deemed “necessary to evaluate the effectiveness of AI technologies in assisting decision making, but also to form a foundational understanding of how people interact with AI to make decisions” (Lai et al., 2023). However, research areas in HCI and related fields that rely heavily on decision-making lack a consensus on what a decision is, and how decisions are a distinct form of human behavior compared to other tasks. Human decisions from prediction and other information displays are often evaluated in studies with the goal of identifying forms of bias: ways in which study participants’ behavior deviates from expectations of good decision-making. But reviewing the range of tasks that get filed under decision-making, it is difficult to ascertain what common task structure makes these studies about decision-making, much less what properties a study needs to have in order for it to be capable of evaluating human decisions, rather than simply describing them.
@@
As examples of the inconsistency and underdefinition of good decision-making in the HCI literature, colloquial definitions describe a decision task as a choice between alternatives, or a choice where the stakes are high (such as deciding which of multiple hurricane forecasts is more accurate (Padilla et al., 2015), or whether to screen child welfare cases (Zytek et al., 2021)). Neither isolates a clear set of assumptions that must hold for a judgment to qualify as a decision. Other accounts cast decision-making as a higher-level task akin to sensemaking (Dimara and Stasko, 2021). For some, the point of studying human decisions is to compare them to a definition of idealized behavior. Sometimes researchers define this standard in a way that is common across studies (e.g., a better forecast comes closer to the true probability of the event (Green and Chen, 2019; Fogliato et al., 2021)). Other times, however, different definitions of the same concept are used: e.g., appropriate reliance as agreeing with an AI when it is correct and disagreeing when it’s not (Bansal et al., 2021; Buçinca et al., 2021), versus providing a higher subjective rating of trust when the model is correct (Bussone et al., 2015), versus choosing the better of the human versus the AI prediction to bet on for a future task (Dietvorst et al., 2015). Sometimes the point is to study seemingly “subjective decisions,” like how people react to music recommendations (Kulesza et al., 2012), or decide whether a house seems to match its valuation (Prabhudesai et al., 2023), where it can seem unclear whether normative interpretations of task performance can be applied. Not only is a more coherent definition of decision needed, as recent survey papers on decision making in visualization and human-centered AI recommend (Dimara and Stasko, 2021; Lai et al., 2021), but such a definition must clearly define requirements for evaluating decisions by identifying forms of bias or performance loss in human decision-making.
This paper contributes a definition of a decision problem and associated guidelines for practice for the purposes of evaluating human decisions from information displays.
We show how a widely applicable framework for studying decision-making can be snthesized from foundations of statistical decision theory and expected utility theory (Savage, 1972; Steele and Stefánsson, 2015; Von Neumann and Morgenstern, 2007), which rigorously motivate what it means to make good decisions under uncertainty, and information economics, which formalizes information structures that capture the relevant information in a set of signals for a decision problem (Bergemann and Morris, 2019). We argue that for a decision problem to be well-defined, it must be possible to rigorously define a normative (best response) decision to the information provided to the study participants. We motivate this criterion by showing how performance loss can only be conceived of when a study participant could in theory identify how to optimize their behavior from the information they are provided. To assess how close recent practice comes to achieving this ideal, we apply our definition to 46 recent studies on AI-assisted decision-making, finding that while 39 of the studies draw conclusions about shortcomings in human decisions with prediction displays, only 10 (26%) of these provided participants a well-defined decision problem in at least one treatment condition.
This paper is organized as follows. First, we define a decision problem, normative decision, and sources of performance loss given human decision-makers. Next, we discuss considerations of scope and important study design decisions in applying this framework in practice. We then present two example study designs in detail, and describe how they could be improved to align with the framework we present. We apply the framework to a sample of studies on AI-assisted decisions, illustrating the gap between the ideal it presents and current practice. We conclude with recommendations for experimenters and a discussion of common misunderstandings of the approach we present.
@@ 2. Defining a Decision Problem
We define a decision problem and the normative decision given such a problem and enumerate possible sources of loss.
@@ 2.1. Decision Problem
The data-generating model is a distribution over values of the uncertain state and signals that provide information to the agent about the state. The data-generating model associates the state with the information that the agent is presented with (the signal) in order to inform their decision; for example a model prediction or visualization of relevant data. Since the signal is correlated with the state, if the participant understands it in the context of the decision problem, they can improve their performance compared to a case where they do not have access to the signal. More precisely, the data-generating model defines a joint distribution (or information structure) over signals and states . This joint distribution assigns to each realization a probability, denoted .
@@
When we present a person (hereafter agent) with a decision problem, we give them a choice of response (action) from some set of possible responses (action space) . The action space can be identical to the state space (as in AI-assisted diagnosis (e.g., (Alur et al., 2024)) or can be defined on a different space. In decision tasks corresponding to prediction tasks, (e.g., “What’s the probability the patient will be readmitted to the hospital?”), the action space is a probabilistic belief over the state space, i.e., . The agent faced with a decision task must consider the probability of different states induced by the signal, and choose the best action under the scoring rule, a function that assigns a score representing the quality of the action for the realized state: .
Under this definition, the formalization of the joint distribution enables a definition of normative behavior. Summarizing information about , (as well as ) provides a basis for the experimenter to establish expectations under optimal use of the information.
@@ 2.2. Normative Behavior
2.2.1. Rational Belief Formation
To define the normative action for a decision problem, we must first characterize how we would ideally expect the agent to form the posterior beliefs about the state given the signal . A theory of optimal belief construction starts by assuming that the ideal (i.e., rational) agent perfectly perceives . Once perceived, how does inform the agent’s beliefs about ? Generally, we assume that the agent has some baseline set of beliefs about that they could use to choose an action in the absence of . Encountering leads the agent to Bayesian update their beliefs.
where is a normalizing factor.
The maximally informative posterior beliefs that any agent can arrive at given the decision problem is:
Note that the above definition differs from the posterior definition in Equation 1 only by explicitly assuming that the agent knows . It is sometimes unreasonable to expect that can be communicated to any agent, rational or not, without incurring some error. For example, in many studies that involve machine learning predictions based on many features and possibly many states (e.g., multi-label classification), communicating the joint distribution over signals and states without loss of information is impractical. Under such conditions, we expect the rational agent to arrive at posterior beliefs defined by:
where represents the approximation of that a rational agent arrives at given the information about employed by the study, such as a sample of labeled instances presented to approximately convey .
2.2.2. Defining the Normative Action
Given the posterior beliefs as defined in Equation 2 or 3, we use Equation 5 to identify the action that a fully rational agent would choose in order to maximize their expected utility under . This is the normative decision.
@@ 2.3. Example
To make these concepts concrete, imagine a study of AI-advised decisions in which people make decisions about whether to book a flight upon seeing features of the flight. The state is whether the flight will be delayed beyond some nominal number of minutes, which can take the value 0 for no or 1 for yes (). The action space presented to the agent is whether or not to book the flight ().
The data-generating model is the joint distribution over the state (delay or not) and signals comprised of flight feature values and AI prediction information. Assume that participants are presented with a signal in the form of a set of feature values describing a flight (origin, destination, time, carrier, prior delays in last 30 days, etc.) from which they must decide on an action. Imagine that the signal also includes the prediction from a model that participants are told was trained on a historical dataset of labeled examples.
For a participant to in theory be able to form posterior beliefs about the probability that a flight is delayed after viewing such signals, they must have prior information about the data-generating model. The joint distribution assigns a probability, for each possible value of ({0, 1}), to each possible signal (a combination of feature values and model prediction) used as a trial in the experiment. Rather than attempting to communicate the exact prior to participants, imagine that before beginning the study trials, participants are provided with a set of labeled instances, which defines . If we did not show the agent any feature values at all (nor model predictions), we would expect them to make their choice based on the probability that a randomly drawn flight in the study will be delayed, , representing the projection of to the state.
Imagine the researcher uses a binary scoring rule to map the ground truth realization of the state ( for no delay, delay) and the participant’s selected action from the action space (), say, booking the flight (), to an outcome. For example, the agent might be scored by the following rule, which imposes the greatest penalty on false negatives (i.e., booking the flight when it will be delayed):
@@
The normative response to this decision problem is defined by the action that a rational agent selects after being presented with the problem (including the action space, scoring rule, and approximation of the data-generating model ), viewing the signal, updating their beliefs, and selecting the score-maximizing action. Figure 1 illustrates this example.